In a weightlifting competition, the amount that the competitors can lift is normally distributed with mean = 196 kg and standard deviation = 11 kg. only the top 20% of all competitors will be able to advance to the next phase of the competition. What amount must a competitor lift in order to move into the next phase of the competition?


1
Expert's answer
2021-02-24T07:10:34-0500

Given,

The mean lifted weight "(w) =196 kg"

Standard deviation "=11kg"

The competitors would have =(100-20)=80%

The z-score at 80 percentile (z)=0.84

"0.84=\frac{x-196}{11}"

"\Rightarrow 11\times0.84=x-196"

"x=205.26kg"